# ChartOptimiser: Task-driven Optimisation of Chart Designs

### v1

@@ 4.3. Parameter Analysis
We further investigated the optimised parameters in the design space and compared our ChartOptimiser with both human-designed charts(Masry et al., 2022) and LQ2-optimised designs (Wu et al., 2021) on the 300 charts introduced in Sec. 4.1. We selected aspect ratio and bar width as the parameters for analysis, as they are the most general design factors when humans create a bar chart (Wu et al., 2021). Figure 7 illustrates the relationship between bar widths and aspect ratios across the 300 charts. In the human-designed charts, the most preferred region is an aspect ratio of 1.6 and 2.0 when the bar width is between 140 and 160 pixels. Our ChartOptimiser closely aligned with these human-designed results, with the most frequent region having an aspect ratio of 1.6 and 2.0 when the bar width is between 120 and 140 pixels. However, the baseline approach LQ2 yielded substantially different results from those of humans, with an aspect ratio of 0.4 and 0.8 when the bar width is between 80 and 100 pixels. In addition, we find that ChartOptimiser covers the parameter space more comprehensively than human and LQ2. This shows that our design can be adapted to diverse design combinations.
@@ 5.1. On ChartOptimiser
Furthermore, the design space of ChartOptimiser allows for a wide range of design possibilities, where font sizes, aspect ratio, bar width, and bar colour allow continuous choices. In stark contrast, previous chart optimisation works all limited their design space to discrete values (Micallef et al., 2017; Wu et al., 2021). Despite the expanded design space in ChartOptimiser, Figure 7 shows that the parameter selections closely align with human-generated designs. By efficiently exploring this space through Bayesian Optimisation, ChartOptimiser consistently achieves highly effective chart designs. This adaptability makes it a robust tool for specific applications, such as chart designing for accessibility or content localisation.
@@ 5.3. On Applications
Data Visualisation Literacy is the ability to interpret visual patterns in the visual domain as properties in the data domain (Boy et al., 2014). Users with low Data Visualisation Literacy might need help with complex data interpretation tasks, like comparison or computing derived values. ChartOptimiser has strong competence in optimising charts for those complex tasks, making the charts more engaging. For users with low visual literacy, complex visualisations can be overwhelming. To mitigate this, ChartOptimiser could prioritise lower information density by increasing the white space ratio, thus avoiding clutter and presenting only essential information. This would help users focus on the most critical aspects of the data. The colour design space of ChartOptimiser could be simplified to ensure accessibility. Prioritising contrasting and accessible colour schemes would facilitate quicker comprehension for users with low visual literacy (Börner et al., 2019). Moreover, ensuring that charts adhere to colour guidelines for colour blindness, like avoiding red-green contrasts, makes visualisations more inclusive and comprehensible to a broader audience.
@@
Journalists and media outlets frequently use visualisations to convey data to the public. Using ChartOptimiser can enhance these visualisations by adaptively emphasising different data points for different audiences, a process known as content localisation (Kee and Yazdanifard, 2015). Figure 8 visualises the percentage of people accessing a personal computer in different countries. For instance, the top figure fits the user located in Jordan, as Jordan is highlighted in the chart. The bottom figure fits the user with interest (e.g. travel plans) in Kenya and India. ChartOptimiser ensures that the charts are visually appealing and optimises them for relevance to specific audiences based on their demographic information (user profile), making it easier for readers to extract key insights and increase engagement with localised stories. Nevertheless, more engineering work is needed given localisation differences, such as reversing the order of axis labels given a language reading from right to left, and choosing culturally plausible colour schemes.
@@ 5.4. Limitations and Future Work
Currently, ChartOptimiser focuses on bar charts, which do not cover the full spectrum of chart types used in practice. Expanding the design space to include other chart types, such as line or scatter plots, could increase the versatility of ChartOptimiser. Another avenue for future work is incorporating dynamic, multi-task visualisations, where charts are optimised to simultaneously support multiple user tasks. In future work, as data storytelling continues to play a pivotal role in data visualisation (Shi et al., 2020; Shao et al., 2024), integrating narrative-driven elements into the automatic chart design process presents a promising direction (Deng et al., 2022). Extending ChartOptimiser to incorporate storytelling could support a wider range of applications, particularly in dashboard design and data communication, where contextualised narratives are essential. Furthermore, ChartOptimiser holds the potential to be integrated into a “ChartGPT” system (Tian et al., 2024), where optimised charts are dynamically generated as responses to user queries, offering tailored data insights on demand. This direction would enhance interactive data exploration and make data insights more accessible through conversational interfaces.
@@ 6. Conclusion
In this work, we introduced ChartOptimiser– the first method for task-driven optimisation of information visualisations. At its core is a Bayesian approach that optimises chart designs with respect to a novel objective function integrating four perceptual metrics – visual saliency, text legibility, colour preference, and white space ratio. Through empirical evaluation for common analytical tasks, we showed that our method can effectively optimise bar chart designs to enhance chart clarity, aesthetics, and task-solving ease. ChartOptimiser demonstrated significant performance improvements over human-designed and baseline charts, especially in more complex tasks such as comparing or computing derived values.

### v2

@@ 5.3. Parameter Analysis
We further investigated the optimised parameters in the design space and compared our ChartOptimiser with both human-designed charts(Masry et al., 2022) and LQ2-optimised designs (Wu et al., 2021) on the 300 charts introduced in Sec. 4.1. We selected aspect ratio and bar width as the parameters for analysis, as they are the most general design factors when humans create a bar chart (Wu et al., 2021). Figure 7 illustrates the relationship between bar widths and aspect ratios across the 300 charts. In the human-designed charts, the most preferred aspect ratio is between 1.6 and 2.0, with a bar width of between 140 and 160 pixels. Our ChartOptimiser closely aligned with these human-designed results, with the most frequent region having an aspect ratio of 1.6 and 2.0 when the bar width is between 120 and 140 pixels. However, the baseline approach LQ2 yielded substantially different results from those of humans, with an aspect ratio of 0.4 and 0.8 when the bar width is between 80 and 100 pixels. In addition, we find that ChartOptimiser covers the parameter space more comprehensively than human and LQ2. This shows that our design can be adapted to diverse design combinations.
@@
5.4. Other Chart Types
With minor modifications, ChartOptimiser could be adapted to other chart types. For instance, Micallef et al. (2017) employed exhaustive search with a perceptual cost function to optimise scatterplots. In this context, exhaustive search could be replaced with Bayesian optimisation to more efficiently explore the continuous design space, such as the colour and opacity of data points. Here, we further demonstrate the application of ChartOptimiser to multi-column bar charts and circular charts. The following figures generated by GPT-5.1 were all generated by the same prompts as described in subsection 5.1 in a temporary session.
Multi-column Bars
To adapt ChartOptimiser, we expanded the design space by introducing a binary variable that indicates whether the chart is stacked. In addition, group colours were derived from the bar and highlight colour specifications to keep minimal changes to the design space. Weights of the objective function is set to $w_{w}=3$, $w_{c}=2$, $w_{t}=2$, $w_{s}=4$. Sample multi-column bar charts generated from ChartOptimiser and GPT-5.1 were illustrated in Figure 8.
GPT-5.1 exhibits robust capabilities for context-aware design, particularly in encoding group semantics through colour, e.g. mapping Clinton (Democrats) to blue and Trump (Republicans) to red. However, the model frequently generates confabulated designs (Siontis et al., 2024) and deviates from established best practices. As shown in ①, the model incorrectly employed a stacked bar chart where the cumulative percentage exceeded 100%, whereas a non-stacked design would have been topologically appropriate. Furthermore, the model tends to violate the principle of minimalism by introducing visual clutter; ② illustrates the addition of redundant axis labels that diminish the data-ink ratio without providing auxiliary information. Finally, in an attempt to balance the layout ③, GPT-5.1 reduced font sizes to 16–18 px, thereby severely compromising readability. On the contrary, ChartOptimiser consistently yields chart designs that prioritise structural coherence and a high data-ink ratio. Although it lacks the semantic colour intuition of LLMs, it ensures high-level readability through optimised spacing, appropriate font sizes, and maintains data integrity, leading to efficient and effective chart designs.
Figure 8. Comparison of automated design approaches for multi-column bar charts. Top: VegaLite defaults (Source: ChartQA). Middle: ChartOptimiser results. Bottom: GPT-5.1 outputs. Although GPT-5.1 exhibits semantic sensitivity, its outputs often suffer from design confabulation and violations of best design practices.
Sample multi-column bar charts from the ChartQA dataset, and optimised bar charts from ChartOptimiser and GPT-5.1.
Circular Charts (Pie and Doughnut charts)
The pie charts were converted from the VegaLite default bar charts described in subsection 5.1, with sorted data in descending order corresponding to the colour palette Category10. For circular charts, ChartOptimiser has font size (data), highlight data colour, inner radius, and the text position (radius) in the design space. Weights of the objective function is set to $w_{w}=0$, $w_{c}=1$, $w_{t}=1$, $w_{s}=1$, removing the white space ratio from the function. We added one more constraint to make sure the data labels always appear within the doughnut. Sample circular charts generated from ChartOptimiser and GPT-5.1 were illustrated in Figure 9.
GPT-5.1’s output demonstrates several instances of design confabulation. First, the model ① arbitrarily altered the visualisation topology to bar charts while failing to annotate the full dataset. Second, data fidelity was compromised through ② the incorrect aggregation of distinct data entries. Moreover, the model failed to adhere to standard encoding heuristics, evidenced by ③ erroneous data transformations (e.g., converting raw values to percentages where $100\%$ indicates a clear logic failure) and the spatial displacement of textual labels. Additionally, ④ the model removed the legend while assigning an identical green colour to multiple data categories, creating ambiguity. In contrast, ChartOptimiser strictly maintains data fidelity. It enhances visual hierarchy by selectively tuning a single colour to increase the saliency of highlighted data points, while supporting switching types between doughnut and pie charts.
Figure 9. Comparison of automated design approaches for circular charts. Top: VegaLite defaults (Source: ChartQA). Middle: ChartOptimiser results. Bottom: GPT-5.1 outputs.
Sample circular charts from the ChartQA dataset, and optimised bar charts from ChartOptimiser and GPT-5.1.
@@ 6.1. On ChartOptimiser
Advantage of Bayesian optimisation.
We employ Bayesian Optimisation due to its superior performance in black-box optimisation scenarios typical of visualisation design. Unlike evolutionary algorithms and other combinatorial optimisations that rely on exhaustive sampling, BO’s sample-efficient nature makes it highly suitable for chart design tasks. We further distinguish our approach from Reinforcement Learning, which typically treats optimisation as a step-by-step sequence; conversely, ChartOptimiser optimises the design space holistically. Additionally, ChartOptimiser overcomes a primary limitation of contemporary deep learning models: the reliance on curated training data. By utilising explicit, perceptually-grounded objective functions, our method provides a “cold-start” capability that ensures effectiveness even in the absence of large-scale training data.
@@
While this work primarily evaluates ChartOptimiser on bar and circular charts, the underlying framework is designed to be generalisable to a broad spectrum of chart types such as scatter and line plots. This generalisability stems from the usage of a high-level VegaLite grammar for the design space and the reliance on image-based perceptual metrics for the objective function. Moreover, the design space has advanced to continuous choices, such as for colour and aspect ratio. In stark contrast, previous chart optimisation works limited their design space to discrete values (Micallef et al., 2017; Wu et al., 2021). Despite the expanded design space in ChartOptimiser, as shown in Figure 7, the parameter selections closely align with human-generated designs. Despite this vastly expanded search space, our analysis (see Figure 5) confirms that ChartOptimiser efficiently converges on parameter selections that closely align with professional human-generated designs. This capability positions ChartOptimiser as a robust solution for upgrading the suboptimal default settings in tools like Excel and VegaLite, addressing the pervasive default effect (Wu et al., 2021) where users often unknowingly settle for adequate but ineffective visualisations.
Bayesian Optimisation versus Large-Language Models.
While Large-Language Models (LLMs) like ChatGPT have demonstrated impressive performance in generating high-quality images, the results in chart design highlight critical issues in fidelity, efficiency, and effectiveness. First, the most significant limitation of LLMs lies in their probabilistic nature. As observed in qualitative results (see Figure 8), LLMs prioritise good-looking designs over data fidelity, leading to confabulations where the generated chart misrepresents the underlying data values. In contrast, ChartOptimiser operates within a strictly defined parameter space. By treating chart design as a constrained optimisation problem rather than a token prediction task, our approach guarantees data integrity. Second, LLMs could offer a zero-shot and quick solution, but achieving high-quality results with LLMs often necessitates expert-level prompt engineering to correct design flaws, resulting again in a human-in-the-loop process. It is a chicken-and-egg problem that users require to have design knowledge to make LLMs generate satisfying chart designs. Third, while LLMs rely on learned correlations from training data to approximate plausible design, they lack an explicit understanding of human visual perception. This results in charts that may be aesthetically pleasing but functionally deficient for specific tasks. ChartOptimiser bridges this gap by directly maximising a rigorous objective function grounded in perceptual metrics (saliency, colour preference, text legibility, and visual density). This ensures that the resulting designs are not just valid code, but are empirically effective at guiding user attention to the most relevant information for their specific analytical goals.
We believe that ChartOptimiser offers a valuable complement to LLMs, specifically thanks to its fidelitous, effective, and efficient performance. First, the BO process has eliminated human-driven iteration (prompt engineering) and is more efficient compared to GPTs. Second, we noticed that GPT modified data entries in some cases, which raises an issue regarding the validity of the results.
@@ 6.3. On Applications
Data Visualisation Literacy is the ability to interpret visual patterns in the visual domain as properties in the data domain (Boy et al., 2014). Users with low Data Visualisation Literacy may need help with complex data interpretation tasks, such as comparison or computing derived values. ChartOptimiser has strong competence in optimising charts for those complex tasks, making the charts more engaging. For users with low visual literacy, complex visualisations can be overwhelming. To mitigate this, ChartOptimiser could prioritise lower information density by increasing the white space ratio, thus avoiding clutter and presenting only essential information. This would help users focus on the most critical aspects of the data. The colour design space of ChartOptimiser could be simplified to ensure accessibility. Prioritising contrasting and accessible colour schemes would facilitate quicker comprehension for users with low visual literacy (Börner et al., 2019). Moreover, ensuring that charts adhere to colour guidelines for colour blindness, like avoiding red-green contrasts, makes visualisations more inclusive and comprehensible to a broader audience.
@@
Journalists and media outlets frequently use visualisations to convey data to the public. Using ChartOptimiser can enhance these visualisations by adaptively emphasising different data points for different audiences, a process known as content localisation (Kee and Yazdanifard, 2015). Figure 10 visualises the percentage of people accessing a personal computer in different countries. For instance, the top figure is relevant for users located in Jordan, as Jordan is highlighted in the chart. The bottom figure fits the user with interest (e.g. travel plans) in Kenya and India. ChartOptimiser ensures that the charts are visually appealing and optimises them for relevance to specific audiences based on their demographic information (user profile), making it easier for readers to extract key insights and increase engagement with localised stories. Nevertheless, more engineering work is needed given localisation differences, such as reversing the order of axis labels given a language reading from right to left, and choosing culturally plausible colour schemes.
@@
Thumbnailing and Grey-scale Charts
Beyond standard displays, ChartOptimiser can be adapted to highly constrained environments, such as thumbnails and print media. In UI applications like dashboards (Deng et al., 2022; Jiao et al., 2010), size reduction necessitates prioritising immediate recognisability over fine-grained detail. ChartOptimiser could automatically emphasise critical data features while eliding illegible elements like legends or minor axis labels. Similarly, adapting visualisations for greyscale print media (e.g. newspapers) removes colour as an encoding channel. In such scenarios, our system could be reconfigured to compensate for the loss of colour by optimising alternative parameters – such as aspect ratio, bar width, or hatching – to maintain structural coherence and communicative intent under technical limitations.
@@ 6.4. Limitations and Future Work
A significant limitation of the current ChartOptimiser implementation is the requirement for manual specification of the design space and optimisation goal tailored to specific chart types. This manual intervention restricts its practical ability. To enhance ChartOptimiser’s versatility, future work should focus on automatically identifying the chart type. This would allow the method to adapt its internal optimisation parameters autonomously, streamlining the generation process. Furthermore, the current method’s chart coverage is constrained. It is challenging for ChartOptimiser to accommodate the full spectrum of chart types used in practice, such as area charts and map visualisations. Leveraging the contextual memory provided by LLMs represents a viable path toward greater sophistication. This integration would enable ChartOptimiser to generate dynamic and interactive charts that are optimised to flexibly support a combination of concurrent user demands.
While the evaluation of ChartOptimiser employed ChatGPT to represent the state-of-the-art in generative design, its application was restricted to prompt-level JSON editing rather than a fully integrated optimisation process in a deep-learning model. This leaves a significant avenue for future exploration involving more systematic applications of Large Language Models. Future work could investigate hybrid architectures that combine the semantic reasoning of multimodal LLMs with the rigorous constraints of our optimisation engine, potentially unlocking new capabilities for end-to-end, model-driven visualisation synthesis.
Another avenue for future work is data storytelling in data visualisation (Shi et al., 2020; Shao et al., 2024), integrating narrative-driven elements into the automatic chart design process presents a promising direction (Deng et al., 2022). Extending ChartOptimiser to incorporate storytelling could support a wider range of applications, particularly in dashboard design and data communication, where contextualised narratives are essential. Furthermore, ChartOptimiser holds the potential to be integrated into a “ChartGPT” system (Tian et al., 2024), where optimised charts are dynamically generated as responses to user queries, offering tailored data insights on demand. This direction would enhance interactive data exploration and make data insights more accessible through conversational interfaces.
@@ 7. Conclusion
In this work, we introduced ChartOptimiser  – the first method for task-driven optimisation of information visualisations. At its core is a Bayesian approach that optimises chart designs with respect to a novel objective function integrating four perceptual metrics – visual saliency, text legibility, colour preference, and white space ratio. Through empirical evaluation of common analytical tasks, we demonstrated that our method can effectively optimise bar chart designs to enhance chart clarity, aesthetics, and ease of task-solving. ChartOptimiser demonstrated significant performance improvements over human-designed and baseline charts, especially in more complex tasks such as comparing or computing derived values.